# Lower Bounds for the Large Deviations of Selberg’s Central Limit Theorem

### v1

@@ 1. Main Results
for some constant $K_{\alpha}$ that depends on $\alpha$ but not $T$, and where $\mathbf{P}$ stands for the uniform probability on $\tau$. This in turn implied another proof for a sharp upper bound for the fractional moments of $\zeta$ between 00 and $4$ as first proved in [HRS19].
@@
The constant $C_{\alpha}(\delta)$ is explicit, see Equation (19). So is $\delta^{\star}$, as given in Equation (33). For $\alpha$ tending to $\infty$, $\delta^{\star}(\alpha)$ is of order $\log\alpha$, whereas $\delta^{\star}(\alpha)$ remains bounded away from 00 as $\alpha\to 0$. (For technical reason, $\delta$ is redefined in (13), but this does not change the asymptotic behavior.) Our method of proof does not seem to apply at $\delta=0$, although it does imply sharp lower bound on the moments on the critical line. More precisely, it was conjectured by Keating & Snaith [KS00] that for all $k\geq 0$
@@
Sharp lower bounds of this kind were proved in [HS22] and [RS13]. The constant obtained by Heap & Sound is better for large $k$ as it behaves like $k^{-2k^{2}}$. For $k$ approaching 00, their constant tends to 00 whereas the one in Corollary 1 is bounded away from it, as it should. Although, as mentioned in their paper, a modification of their argument would lead to the correct behavior for small $k$. Note that the constant $c_{k}$ includes explicitly the arithmetic part $a_{k}$, and this is in fact necessary in our argument. However, this does not change the asymptotic behavior for large $k$, since as shown in Conrey & Gonek [CG01], $\log a_{k}\sim-k^{2}\log\log k$ as $k\to\infty$. The correct asymptotic behavior of $f_{k}$ is $\log f_{k}\sim-k^{2}\log k$. It is unclear to us at this time how to improve the constant $c_{k}$, though a better approximation of indicator functions in terms of polynomials seems necessary, cf. Lemma 2 and the discussion following Equation (19). It is important to note that the same constant $C_{k}$ as for the moments should appear in the large deviation of $\log|\zeta|$. Indeed, it was conjectured by Radziwiłł that for any $\alpha>0$ and $V\sim\alpha\log\log T$ as $T\to\infty$,
@@ 2.3. Proof of Theorem 1
where the upper and lower barriers $L_{\ell}(z),U_{\ell}(z)$ are defined and discussed below in (25), and $S_{0}=-\frac{1}{2}\log|\mathcal{M}_{0}|$, corresponding to the definition (9) and the choice in the initial good event $G_{0}$. (This event is chosen since $\mathcal{M}_{0}$ is a short Dirichlet polynomial, cf. (43).) We take
@@
where $\log_{\ell}$ stands for the logarithm iterated $\ell$ times. In each case, the slope $m(z)$ will be evaluated at the (random) initial position, i.e. $S_{0}$ or $\mathcal{Z}_{0}$. Taking the barriers randomly is necessitated by separating out the contribution from the small primes. On first reading, one can think of the slope $m(z)\approx\alpha$ (since typically $z$ is like $\alpha\log\log T_{0}$ in each case). However, we need to take in to account the variance coming from the initial random sums which informs the slope depending on the random position of $\mathcal{Z}_{0}$ or $S_{0}$, see in particular the remark following (36). See also Figure 1, where two different initial values $\mathcal{Z}_{0}$ and $\mathcal{Z}^{\prime}_{0}$ both reach level $V$ but elicits two different slopes.
@@ 3.2. Proof of Proposition 2
For the $u_{\ell}$, $1\leq\ell\leq\mathcal{L}$, we ask that the $u$’s satisfy the “barrier” condition,
(45) $L_{k}(u_{0})\leq u_{0}+\sum_{1\leq k\leq\ell}u_{k}\leq U_{k}(u_{0})\text{ for $1\leq\ell\leq\mathcal{L}$.}$
In the following, to ease notation, we write $L_{\ell}(u_{0})=L_{\ell}$. The condition (45) implies the following bounds on the $u_{\ell}$’s, $1\leq\ell\leq\mathcal{L}$:
$u_{\ell}$ $\leq U_{\ell}-L_{\ell-1}=(\mathfrak{s}\,m(u_{0})+\mathcal{C})\log_{\ell-1}t-(\mathfrak{s}\,m(u_{0})-\mathcal{B})\log_{\ell}t,$
$u_{\ell}$ $\geq L_{\ell}-U_{\ell-1}=(\mathfrak{s}\,m(u_{0})-\mathcal{B})\log_{\ell-1}t-(\mathfrak{s}\,m(u_{0})+\mathcal{C})\log_{\ell}t.$
@@
(46) $|u_{\ell}|\leq 2\alpha\mathfrak{s}\log_{\ell-1}t,\quad 1\leq\ell\leq\mathcal{L}.$
This is the important input for the choice of parameters. From this, we take
$\Delta_{\ell}=10\alpha\mathfrak{s}\log_{\ell-1}t,\quad 1\leq\ell\leq\mathcal{L}.$
@@
For $\ell=0$, keeping in line with the choice in (43), we define
@@
$\mathbf{1}(Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}])$ $\geq\mathbf{1}_{[0,\Delta_{\ell}^{-1}]}(Y_{\ell}-u_{\ell})\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})$
$\geq|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell}),$
@@
$\mathbf{1}(Y_{0}\in[u_{0},u_{0}+\Delta_{0}^{-1}])\geq|\mathcal{D}_{0}^{-}(Y_{0}-u_{0})-e^{-\Delta_{0}^{a-2}}|^{2}.$
@@
$\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
(56) $=\sum_{J\subseteq\{1,\dots,\mathcal{L}\}}(-1)^{|J|}\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{j\in J}\mathbf{1}(|Y_{j}-u_{j}|>X_{j})\right].$
@@
$\mathbf{E}\left[\prod_{\ell}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]=(1+\operatorname{O}(T^{-1/100}))\prod_{\ell}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right].$
Each factor in the product is easily bounded below by
(57) $\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]\geq\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{(}1-\frac{2e^{-\Delta_{\ell}^{a-2}}}{[\mathbf{E}[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}]]^{1/2}}\Big{)}.$
@@
$\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]$ $\geq\Big{(}1-e^{-\Delta_{\ell}^{a-2}/2}\Big{)}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$
@@
(64) $\mathbf{E}\left[|\mathcal{D}_{1}^{-}(\mathcal{Y}_{1}-u_{1})-e^{-\Delta_{1}^{a-2}}|^{2}\right]\gg\mathbf{P}(\mathcal{N}_{1}\in[u_{1},u_{1}+\Delta_{1}^{-1}]).$
We now establish a similar bound for $\mathcal{Y}_{0}$. Lemma 2 implies that
@@
(66) $\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})-e^{-\Delta_{0}^{a-2}}|^{2}\right]\gg\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})|^{2}\right]\gg\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}]).$
@@
It remains to estimate the summands in (56) with $J\neq\emptyset$. For each $J$, we can bound the indicator functions to get
(68) $\leq\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{j\in J}X_{j}^{-2q_{j}}|Y_{j}-u_{j}|^{2q_{j}}\right],$
for an appropriate choice of $q_{j}$, $j\in J$. The length of the polynomial in the expectation is then at most
@@
The mean-value theorem of Dirichlet polynomial can be applied if $q_{j}$ is not too large, for example $q_{j}\leq X_{j}\Delta_{j}$ would do. Passing to the random model, Equation (68) is
(70) $=\prod_{\ell\in J^{c}\cup\{0\}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]\prod_{j\in J}X^{-2q_{j}}\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{2}|\mathcal{Y}_{j}-u_{j}|^{2q_{j}}\right].$
The expectation on the right is for each $j$
(71) $\leq\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{4}\right]^{1/2}\cdot\mathbf{E}[|\mathcal{Y}_{j}-u_{j}|^{4q_{j}}]^{1/2}.$
We evaluate both terms separately. By Equations (62) for $j>1$ and Equations (64), (65) for $j=1$ and $j=0$, we have
$\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}\right]^{1/2}\geq\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{2}\right]\gg e^{-\Delta_{j}^{2}}.$
This implies
$\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{4}\right]^{1/2}$ $\leq 2\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}\right]^{1/2}+2e^{-2\Delta_{j}^{a-2}}$
$\leq 4\mathbf{E}[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}]^{1/2}.$
The fourth moment for $j\geq 1$ is estimated using the bounds on the coefficients in Lemma 2 and the elementary estimates on the random model, see Equation (113) in Lemma 10,
@@
$\leq\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\prod_{j\in J}\frac{\exp(50\Delta_{j}^{4a}-2X_{j})}{\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{2}\right]}.$
This leads to the choice
@@
Together with the bound (62), this shows that the above is $\leq\prod_{j\in J}e^{-\Delta_{j}^{4a}}$. Combining this with the dominant term (63), we obtain
$\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{1}(|\mathcal{Y}_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
$\geq\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{)}\Big{(}1-e^{-\Delta^{a}_{\mathcal{L}}/2}-\sum_{J\subset\{1,\dots,\mathcal{L}\},J\neq\emptyset}\prod_{j\in J}e^{-\Delta_{j}^{4a}}\Big{)}$
$=\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{)}\Big{(}2-e^{-\Delta_{\mathcal{L}}^{a}/2}-\prod_{j\leq\mathcal{L}}(1+e^{-\Delta_{j}^{4a}})\Big{)}.$
This is $\gg\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$ for $\Delta_{\mathcal{L}}$ large enough. The final result is Equation (67). To re-express the events in terms of the random model, we use the inclusion (49) with $\mathcal{Y}_{0}$ and $\mathcal{Z}_{\ell}$. This proves the claim (53).

### v2

@@ 1. Main Results
for some constant $K_{\alpha}$ that depends on $\alpha$ but not $T$, and where $\mathbf{P}$ stands for the uniform probability on $\tau$. This in turn implied another proof for a sharp upper bound for the fractional moments of $\zeta$ between 00 and $4$ as first proved in [HRS19]. The large deviations of the imaginary part of $\log\zeta$ for a wide range values are investigated in [Dob24].
@@
The constant $C_{\alpha}(\delta)$ is explicit, see Equation (19). So is $\delta^{\star}$, as given in Equation (33). For $\alpha$ tending to $\infty$, $\delta^{\star}(\alpha)$ is of order $\log\alpha$, whereas $\delta^{\star}(\alpha)$ remains bounded away from 00 as $\alpha\to 0$. (For technical reasons, $\delta$ is redefined in (13), but this does not change the asymptotic behavior.) Our method of proof does not seem to apply at $\delta=0$, although it does imply sharp lower bound on the moments on the critical line. More precisely, it was conjectured by Keating & Snaith [KS00] that for all $k\geq 0$
@@
Sharp lower bounds of this kind were proved in [HS22] and [RS13]. The constant obtained by Heap & Soundararajan is better for large $k$ as it behaves like $k^{-2k^{2}}$. For $k$ approaching 00, their constant tends to 00 whereas the one in Corollary 1 is bounded away from it, as it should. Although, as mentioned in their paper, a modification of their argument would lead to the correct behavior for small $k$. Note that the constant $c_{k}$ includes explicitly the arithmetic part $a_{k}$, and this is in fact necessary in our argument. However, this does not change the asymptotic behavior for large $k$, since as shown in Conrey & Gonek [CG01], $\log a_{k}\sim-k^{2}\log\log k$ as $k\to\infty$. The correct asymptotic behavior of $f_{k}$ is $\log f_{k}\sim-k^{2}\log k$. It is unclear to us at this time how to improve the constant $c_{k}$, though a better approximation of indicator functions in terms of polynomials seems necessary, cf. Lemma 2 and the discussion following Equation (19). It is important to note that the same constant $C_{k}$ as for the moments should appear in the large deviation of $\log|\zeta|$. Indeed, it was conjectured by Radziwiłł that for any $\alpha>0$ and $V\sim\alpha\log\log T$ as $T\to\infty$,
@@ 2.3. Proof of Theorem 1
where the upper and lower barriers $L_{\ell}(z),U_{\ell}(z)$ are defined and discussed below in (25), and $S_{0}=-\log|\mathcal{M}_{0}|$, following the definition (9). The initial good event $G_{0}$ is defined in terms of $\mathcal{M}_{0}$, since it is a short Dirichlet polynomial, cf. (43). We take
@@
where $\log_{\ell}$ stands for the logarithm iterated $\ell$ times. In each case, the slope $m(z)$ will be evaluated at the (random) initial position, i.e., $S_{0}$ or $\mathcal{Z}_{0}$. Taking the barriers randomly is necessitated by separating out the contribution from the small primes. On first reading, one can think of the slope $m(z)\approx\alpha$ (since typically $z$ is like $\alpha\log\log T_{0}$ in each case). However, we need to take in to account the variance coming from the initial random sums which informs the slope depending on the random position of $\mathcal{Z}_{0}$ or $S_{0}$, see in particular the remark following (36). See also Figure 1, where two different initial values $\mathcal{Z}_{0}$ and $\mathcal{Z}^{\prime}_{0}$ both reach level $V$ but elicits two different slopes.
@@ 3.2. Proof of Proposition 2
For the $u_{\ell}$, $1\leq\ell\leq\mathcal{L}$, and $s_{0}=-\frac{1}{2}\log u_{0}$, we ask that the $u$’s satisfy the “barrier” condition,
(45) $L_{\ell}(s_{0})\leq s_{0}+\sum_{1\leq k\leq\ell}u_{k}\leq U_{\ell}(s_{0}).$
In the preceding and following, we write $L_{\ell}=L_{\ell}(s_{0})$. The condition (45) implies that for $1\leq\ell\leq\mathcal{L}$:
$L_{\ell}-U_{\ell-1}\leq u_{\ell}\leq U_{\ell}-L_{\ell-1}.$
For $\ell\geq 2$ this yields the bounds
$u_{\ell}$ $\leq U_{\ell}-L_{\ell-1}=(\mathfrak{s}\,m(s_{0})+\mathcal{C})\log_{\ell-1}t-(\mathfrak{s}\,m(s_{0})-\mathcal{B})\log_{\ell}t,$
$u_{\ell}$ $\geq L_{\ell}-U_{\ell-1}=(\mathfrak{s}\,m(s_{0})-\mathcal{B})\log_{\ell-1}t-(\mathfrak{s}\,m(s_{0})+\mathcal{C})\log_{\ell}t.$
@@
(46) $|u_{\ell}|\leq 2\alpha\mathfrak{s}\log_{\ell-1}t,\quad 2\leq\ell\leq\mathcal{L}$
and $|u_{1}|\leq 2\alpha t_{1}$. This is the important input for the choice of parameters. From this, we take
$\Delta_{\ell}=10\alpha\mathfrak{s}\log_{\ell-1}t,\quad 2\leq\ell\leq\mathcal{L};\qquad\Delta_{\ell}=10\alpha t_{1}.$
@@
with $\mathcal{Y}_{1}=\mathcal{S}_{1}$. For $\ell=0$, keeping in line with the choice in (43), we define
@@
(54) $\mathbf{1}(Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}])$ $\geq\mathbf{1}_{[0,\Delta_{\ell}^{-1}]}(Y_{\ell}-u_{\ell})\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})$
$\geq\left(|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell}),$
@@
(55) $\mathbf{1}(Y_{0}\in[u_{0},u_{0}+\Delta_{0}^{-1}])\geq|\mathcal{D}_{0}^{-}(Y_{0}-u_{0})|^{2}-e^{-\Delta_{0}^{a-2}}.$
@@
$\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}\left(|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
(58) $=\sum_{J\subseteq\{1,\dots,\mathcal{L}\}}(-1)^{|J|}\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}\left(|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\prod_{j\in J}\mathbf{1}(|Y_{j}-u_{j}|>X_{j})\right].$
@@
(59) $\mathbf{E}\left[\prod_{\ell}\left(|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\right]=(1+\operatorname{O}(T^{-1/100}))\prod_{\ell}\mathbf{E}\left[\left(|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\right].$
@@
$\mathbf{E}\left[\left(|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}-e^{-\Delta_{\ell}^{a-2}}\right)\right]$ $\geq\Big{(}1-e^{-\Delta_{\ell}^{a-2}+\Delta_{\ell}^{2}}\Big{)}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$
@@
(66) $\mathbf{E}\left[|\mathcal{D}_{1}^{-}(\mathcal{Y}_{1}-u_{1})|^{2}-e^{-\Delta_{1}^{a-2}}\right]\gg\mathbf{P}(\mathcal{N}_{1}\in[u_{1},u_{1}+\Delta_{1}^{-1}])$
where $\mathcal{N}_{1}$ is a centred Gaussian with variance $v_{1}^{2}$. We now establish a similar bound for $\mathcal{Y}_{0}$. Lemma 2 implies that
@@
(68) $\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})|^{2}-e^{-\Delta_{0}^{a-2}}\right]\gg\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})|^{2}\right]\gg\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}]).$
@@
It remains to estimate the summands in (58) with $J\neq\emptyset$. For each $J$, we get that each summand is bounded by (using that $\mathbf{1}\{a<b\}\leq(b/a)^{2q}$)
(70) $\leq\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}(d_{\ell}(Y_{\ell})-\varepsilon_{\ell})\prod_{j\in J}X_{j}^{-2q_{j}}|Y_{j}-u_{j}|^{2q_{j}}\right],$
for an appropriate choice of $q_{j}$, $j\in J$, and writing $d_{\ell}(y)=|D_{\ell}^{-}(y-u_{\ell})|^{2}$, $\varepsilon_{\ell}=e^{-\Delta_{\ell}^{a-2}}$. We will show that this is subdominant to the contribution from $J=\emptyset$. Expanding the first factor we get
(71) $\sum_{K\subseteq\{0,\dots,\mathcal{L}\}}\prod_{k\in K^{c}}\varepsilon_{k}\mathbb{E}\left[\prod_{\ell\in K}d_{\ell}(Y_{\ell})\prod_{j\in J}X_{j}^{-2q_{j}}|Y_{j}-u_{j}|^{2q_{j}}\right].$
The length of the polynomial in the expectation is then at most
@@
The mean-value theorem of Dirichlet polynomials can be applied if $q_{j}$ is not too large, for example $q_{j}\leq X_{j}\Delta_{j}$ suffices. Passing to the random model, Equation (70) is
(73) $\leq\sum_{K\subseteq\{0,\dots,\mathcal{L}\}}\prod_{k\in K^{c}}\varepsilon_{k}\prod_{j\in J}X_{j}^{-2q_{j}}\prod_{\ell\in K\cap J^{c}}\mathbb{E}\left[d_{\ell}(\mathcal{Y}_{\ell})\right]\prod_{j\in K\cap J}\mathbb{E}\left[d_{j}(\mathcal{Y}_{j})|\mathcal{Y}_{j}-u_{j}|^{2q_{\ell}}\right]\prod_{j\in J\cap K^{c}}\mathbb{E}\left[|\mathcal{Y}_{j}-u_{j}|^{2q_{j}}\right].$
For the term with $j\in K\cap J$, we may split the expectation using
(74) $\ll\mathbb{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}\right]^{1/2}\mathbb{E}\left[\mathcal{Y}_{j}-u_{j}|^{4q_{j}}\right]^{1/2}.$
We now estimate the moments of $\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})$ and $|\mathcal{Y}_{j}-u_{j}|$ separately. The fourth moment for the former, for $j\geq 1$, is estimated using the bounds on the coefficients in Lemma 2 and the elementary estimates on the random model, see Equation (116) in Lemma 10,
@@
$\leq\sum_{K\subseteq\{0,\dots,\mathcal{L}\}}\prod_{k\in K^{c}}\varepsilon_{k}\prod_{j\in J}e^{50\Delta_{j}^{4a}-2X_{j}}\prod_{j\in K\cap J^{c}}\mathbb{E}\left[d_{j}(\mathcal{Y}_{j})\right]$
$=\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbb{E}\left[d_{\ell}(\mathcal{Y}_{\ell})\right]\Big{)}\sum_{K\subseteq\{0,\dots,\mathcal{L}\}}\prod_{j\in K\cap J}\frac{e^{50\Delta_{j}^{4a}-2X_{j}}}{\mathbb{E}[d_{j}(\mathcal{Y}_{j})]}\prod_{j\in K^{c}\cap J}\varepsilon_{j}\cdot\frac{e^{50\Delta_{j}^{4a}-2X_{j}}}{\mathbb{E}[d_{j}(\mathcal{Y}_{j})]}\prod_{j\in K^{c}\cap J^{c}}\frac{\varepsilon_{j}}{\mathbb{E}\left[d_{j}(\mathcal{Y}_{j})\right]}$
$=\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbb{E}\left[d_{\ell}(\mathcal{Y}_{\ell})\right]\Big{)}\prod_{j\in J}\frac{e^{50\Delta_{j}^{4a}-2X_{j}}}{\mathbb{E}[d_{j}(\mathcal{Y}_{j})]}\Big{(}\sum_{K\subseteq\{0,\dots,\mathcal{L}\}}\prod_{j\in K^{c}\cap J}\varepsilon_{j}\prod_{j\in K^{c}\cap J^{c}}\frac{\varepsilon_{j}}{\mathbb{E}\left[d_{j}(\mathcal{Y}_{j})\right]}\Big{)}$
$\ll\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbb{E}\left[d_{\ell}(\mathcal{Y}_{\ell})\right]\Big{)}\prod_{j\in J}\frac{e^{50\Delta_{j}^{4a}-2X_{j}}}{\mathbb{E}[d_{j}(\mathcal{Y}_{j})]},$
where we used the bound (64) and the value of $\varepsilon_{j}$ in the last inequality. This leads to the choice
@@
This shows that the terms with $J\neq\emptyset$ are indeed subdominant to (69). The claim (53) follows after summing over ${\mathbf{u}}\in\mathcal{I}$ and appealing to the inclusion (49) with $\mathcal{Y}_{0}$ and $\mathcal{Z}_{\ell}$.